OpenStudy (anonymous):

(2s+5)/(s^2+6s+34) inverse Laplace transform

OpenStudy (anonymous):

I have the answer. thank you very much! :)

OpenStudy (anonymous):

.

OpenStudy (anonymous):

are you transferring it back into function F(t)

OpenStudy (anonymous):

someone help me lol

OpenStudy (anonymous):

(2s+5)/(s^2+6s+9+25) = (2s+6-1)/((s+3)^2+25) = (2(s+3)-1)/((s+3)^2+5^2) = 2[(s+3)/((s+3)^2+5^2) - 1/5 (5 / ((s+3)^2+5^2)) Hence inverse Laplace transform is 2e^(-3t)cos 5t - (1/5) e^(-3t)sin 5t

OpenStudy (anonymous):

Does the answer match?

OpenStudy (anonymous):

\[e^{-3t}(2\cos(5t)-\frac{\sin5t}{5})\]

OpenStudy (anonymous):

cool it matches :)

OpenStudy (anonymous):

are making me doubt my answer a moment and rectify